OpenStudy (anonymous):

Show the product of (4√3-6√2)(3√3+2√2). Please help,thank you

OpenStudy (matt101):

Expand and simplify: (4√3-6√2)(3√3+2√2) = 12(3) + 8√6 - 18√6 - 12(2) = 12 - 10√6